OpenStudy (swissgirl):

Can someone help me differentiate \[\int_0^1e^{-st}tdt + \int_1^{\infty} e^{-st}(2-t)dt\]

OpenStudy (swissgirl):

My answer differs from the book unfortunately -_-

OpenStudy (amistre64):

looks lagrangian

OpenStudy (amistre64):

or is it laplace? i mix those guys up alot

OpenStudy (swissgirl):

Laplace transform

OpenStudy (swissgirl):

But like all you gotta do is integrate

OpenStudy (amistre64):

\[\int_0^1e^{-st}t~dt + \int_1^{\infty} e^{-st}(2-t)~dt\] well, a by parts on the poly usually sheds some light if youve forgotten the tables, and i tend to forget the tables e^-st t e^-st/ -s -1 e^-st/ (-s)^2 0 \[\int_0^1e^{-st}t~dt=-\frac{t}{s}e^{-st}-\frac{1}{(-s)^2}e^{-st}\] at t=0 and t=1

OpenStudy (amistre64):

\[\int_1^{\infty} e^{-st}(2-t)~dt\] (2-t) same run down 1 on this side 0 \[\int_1^{\infty} e^{-st}(2-t)~dt=e^{-st}(-\frac{2-t}{s}+\frac{1}{(-s)^2})\]

OpenStudy (swissgirl):

Ok thanks. Not sure what tables u r referring to though

OpenStudy (amistre64):

laplace tranform tables, they present them after they make you sweat this stuff out.

OpenStudy (swissgirl):

ohhhhhhh Thats what u mean. Ya cant use those I gotta do it by hand :(

OpenStudy (amistre64):

\[-e^{-s}(\frac{1}{s}+\frac{1}{(-s)^2})+e^{0}(\frac{0}{s}+\frac{1}{(-s)^2})\]plus \[\cancel{e^{-\infty}(-\frac{2-(\infty)}{s}+\frac{1}{(-s)^2})}^0-e^{-s}(-\frac{2-1}{s}+\frac{1}{(-s)^2})\]

OpenStudy (swissgirl):

Yaaaa thats the answer in the book not sure where I went wrong THANKS :)

OpenStudy (amistre64):

youre welcome; the error tends to be in the algebra with these things :)

OpenStudy (swissgirl):

Yaaaa and I can never find my mistakes. Thanks :)